# Passage-specific Prompt Tuning for Passage Reranking in Question Answering with Large Language Models

### v1

@@ Abstract.
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@ 2.2. Prompt Tuning
Figure 1. The architecture of the proposed PSPT. The original LLM parameters $\Phi$ (green blocks) are frozen during training, with only $\theta$ parameters (red and yellow blocks) updated.
@@ 4.1. Datasets, Baselines and Evaluation Metrics
Dataset Train Ret.Train Eval Test
Natural Questions 79,168 58,880 8,757 3,610
TriviaQA 78,785 60,413 8,837 11,313
SQuAD 78,713 70,096 8,886 10,570
Following the existing work of DPR (Karpukhin et al., 2020) and aiming for fair comparisons in QA tasks, our study utilized several widely used Open-domain question answering datasets including Natural Questions (NQ) (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), and SQuAD (Rajpurkar et al., 2016). The statistics of our datasets are presented in Table 1.
@@
In our work, we employed Llama-2 (Touvron et al., 2023). Specifically, we utilized the fine-tuned Llama-2-Chat model with 7 billion parameters [^2] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf, which was fine-tuned using publicly available instruction datasets. We employed the top-$k$ Recall (R@$k$) and Hit Rate (H@$k$) to evaluate the reranking performance.
@@
Table 2. Comparative analysis of Top-10 Recall and Hit Rate (R@10 and H@10) across test datasets, highlighting the best results for each retriever in bold. The symbols $\dagger$ and $\ddagger$ indicate statistically significant improvements over basic retrievers and the UPR approach, respectively, determined by two-tailed t-tests with p-values ¡ 0.01.
@@ 4.2. Implementation Details
For the baseline models, we used the base configuration as specified in each respective paper. We implemented PSPT based on the publicly available prompt tuning package PEFT (Mangrulkar et al., 2022). We choose hard prompt initialization $s$ as “please generate question for this passage” with pre-defined soft prompt length $l_{s}=50$. For hyper-parameters, $r=1$ and $\alpha=16$ are selected for learning passage-specific embedding $e_{2}$ in Figure 1. We fine-tuned the PSPT model on Nvidia A100 GPUs, using the bfloat16 (Kalamkar et al., 2019) data type, across different training sample sizes ranging from 320 to 1280 for each dataset. The training involved a batch size of 4 and an in-batch hard negative sample size of 4. The learning rates for yellow blocks and red blocks in Figure 1 are set at 3e-2 and 3e-5, respectively, with each undergoing linear decay. A maximum of 20 fine-tuning epochs were executed, incorporating early stopping based on validation set performance. We have displayed additional results of hyper-parameter tuning in Appendix A.
@@ 4.3. Experimental Results
Figure 2. Performance analysis of key components in PSPT, evaluated using BM25 and MSS-DPR on a sampled NQ Dataset with Recall (R@10).
@@
Table 3. Performance comparison of PSPT modules on BM25, MSS-DPR, evaluated on the sampled NQ Dataset using Recall (R@10) and Hit Rate (H@10), highlighting the best results in bold. Each experiment shows the percentage of trainable parameters relative to Llama-2’s total parameters. FT and LoRA denote fully fine-tuning and LoRA-based tuning of the passage-specific module, respectively.

### v2

@@ Abstract.
journalyear: 2024conference: The Second Workshop on Generative Information Retrieval; July 18, 2024; Washington DC, USbooktitle: The Second Workshop on Generative Information Retrieval (Gen-IR@SIGIR24), July 18, 2024, Washington DC, UScopyright: rightsretained
@@ 2.2. Prompt Tuning
Figure 1. The architecture of PSPT. The original LLM parameters $\Phi$ (green blocks) are frozen during training, with only $\theta$ parameters (red and yellow blocks) updated.
@@ 4.1. Datasets, Baselines and Evaluation Metrics
Following the existing work of DPR (Karpukhin et al., 2020) and aiming for fair comparisons, our study utilized the QA datasets presented in Appendix A.6.
@@
In our work, we utilized the Llama-2-Chat model with 7 billion parameters. We employed the top-$k$ Recall (R@$k$) and Hit Rate (H@$k$) to evaluate the reranking performance.
@@
Table 1. The symbols $\dagger$ and $\ddagger$ indicate statistically significant improvements over basic retrievers and the UPR approach, respectively, determined by t-test with p-values ¡ 0.05.
@@ 4.2. Implementation Details
For the baseline models, we used the base configuration as specified in each respective paper. We implemented PSPT based on the publicly available prompt tuning package PEFT (Mangrulkar et al., 2022). We choose hard prompt initialization $s$ as “please generate question for this passage” with pre-defined soft prompt length $l_{s}=50$. For hyper-parameters, $r=1$ and $\alpha=16$ are selected for learning passage-specific embedding $e_{2}$ in Figure 1. We fine-tuned the PSPT model on Nvidia A100 GPUs, using the bfloat16 (Kalamkar et al., 2019) data type, across different training sample sizes ranging from 320 to 1280 for each dataset. The training involved a batch size of 4 and an in-batch negative sampling. The learning rates for yellow blocks and red blocks in Figure 1 are set at 3e-2 and 3e-5, respectively, with linear decay. We trained PSPT 20 epochs with early stopping. We have displayed additional results of hyper-parameter tuning in Appendix A.
@@ 4.3. Experimental Results
Figure 2. Performance analysis of key components in PSPT on the sampled NQ Dataset.
@@
Table 2. Comparison of PSPT modules on the sampled NQ Dataset. Trainable parameters relative to Llama-2’s total parameters are presented. FT and LoRA denote fully fine-tuning and LoRA-based tuning of the passage-specific module, respectively.
@@ A.5. Different Training Samples
A.6. Data Description
Dataset Train Ret.Train Eval Test
Natural Questions (Kwiatkowski et al., 2019) 79,168 58,880 8,757 3,610
TriviaQA (Joshi et al., 2017) 78,785 60,413 8,837 11,313
SQuAD (Rajpurkar et al., 2016) 78,713 70,096 8,886 10,570